Standard Generators for J_{3}

Ibrahim A. I. Suleiman Robert A. Wilson

Abstract

We present a set of generators for the sporadic group J_{3} designed to allow easily reproducible computations in the group. We also discuss the relationship between J_{3} and J_{3}:2, and the maximal subgroups of these two groups.